Plant scientists developed different varieties of corns that have a rich content of lysine which is a nutritious animal feed. A group of chicks were given this food to test the quality. The distribution of the weight gains (in grams) of these chicks are shown below:


Weight gains (in grams) Frequency

318 - 335 4

336 - 353 5

354 - 371 2

372 - 389 3

390 – 407 2

408 – 425 3

426 - 443 1


Find:

(a) the mean weight gains

(b) the median

(c) the variance for the above frequency intervals

(d) the standard deviation


1
Expert's answer
2021-03-30T08:05:31-0400

Construct a table

"\begin{array}{|c|c|c|c|}\hline
\rm{Intervals} & \rm{Frequencies} & \rm{Midpoints\ of\ intervals}&\rm{Cumulative\ frequencies}\\ \hline
318 - 335& 4& 326.5& 4\\ \hline
336 - 353& 5& 344.5& 9\\ \hline
354 - 371& 2& 362.5& 11\\ \hline
372 - 389& 3& 380.5& 14\\ \hline
390 - 407& 2& 398.5& 16\\ \hline
408 - 425& 3& 416.5& 19\\ \hline
426 - 443& 1& 434.5& 20\\ \hline
\end{array}"


The total number of observations is

"n=4+5+2+3+2+3+1=20"


(a) The mean weight gains is


"\overline{x}=\dfrac{1}{20}\cdot(4\cdot326.5+5\cdot344.5+2\cdot362.5+"


"3\cdot380.5+2\cdot398.5+3\cdot416.5+1\cdot434.5)="

"\dfrac{1}{20}\cdot7376=368.8"



(b) The median

The first interval with cumulative frequence ">\dfrac{n}{2}" is "[354-371]" . Previous interval has the frequency "=9" . The length of any interval is "18" . The total number of observations is "20". So the median is

"Me=354+18\cdot\dfrac{\frac12\cdot20-9}{2}=363"


(c) The variance


"\bold{Var}[X]=\overline{x^2}-\overline{x}^2="


"\dfrac{1}{20}\cdot(4\cdot(326.5)^2+5\cdot(344.5)^2+2\cdot(362.5)^2+"


"3\cdot(380.5)^+2\cdot(398.5)^2+3\cdot(416.5)^2+1\cdot(434.5)^2)-"


"(368.8)^2=" "137,188.75-136,013.44=1,175.31"



(d) The standard deviation


"s=\sqrt{\bold{Var}[X]}=\sqrt{1,175.31}\approx34.28"